# eMoE-Tracker: Environmental MoE-based Transformer for Robust Event-guided Object Tracking

### v2

@@ IV-C Ablation Studies
Effectiveness of eMoE and CRM. To validate the effectiveness of our proposed modules, we perform the ablation study on the VisEvent and COESOT datasets. We implement four comparison experiments inside the network. They are: 1) backbone network 2) backbone network with eMoE; 3) backbone network with eMoE and CRM. The ablation studies can be found in TABLE III. As observed, the best performance happens when we combine all the proposed modules into the backbone network. For the VisEvent dataset, with the incorporation of eMoE and CRM, our method outperforms the backbone model by 7.9% and 6.9% on metrics SR and PR, respectively. Additionally, adding the CRM module to model ①, the SR and PR gain improvements of 2.1% and 0.6%, respectively. The results showcase the effectiveness of the proposed modules.
Analysis on environmental attributes. In the dataset COESOT, there are 17 challenging environmental attributes are annotated to help analyze the performance under different challenging conditions. Here we illustrate the overall performance on the COESOT dataset for the disentangled four challenging attributes: illumination variation, motion blur, scale variance, and full occlusion. In Fig. 7, we can find our proposed model eMoE-Tracker outperforms the backbone model and ViPT [10]. Specifically, it achieves 63.6% in occlusion, 77.9% in illumination variance, 73.5% in motion blur, and 80.7% in scale variance on PR, respectively. The results demonstrate that our algorithm is effective in improving tracking precision and robustness under various challenging scenarios.
Inserted layers of eMoE. We achieve RBG-E tracking under various challenging conditions by injecting eMoE blocks into different layers of the backbone model. It is intuitive to investigate the effect on the number of inserted blocks. Here we set different insert intervals for blocks to insert and the intervals are 1,2 4, 6, and 12. Therefore, the first means that all the layers are fully inserted and the last one only inserts the blocks in the last high-level layer.
@@
Analysis on the number of experts. Due to the complicated environmental attributes, it is worthwhile to consider the impact of the number of experts on tracking performance. Intuitively, more experts are more powerful at addressing complex environments and decomposing them into environmental attributes for easier tracking. However, on one hand, too many experts increase the burden on the model because of increasing parameters, on the other hand, it might result in an overfitting phenomenon. Therefore, we conduct ablation studies on the number of experts, and the results can be found in TABLE IV.
Analysis on the model complexity. As we mentioned previously, trackers with the two-stream structure, e.g.siamese-based trackers, suffer from model complexity due to the high demand for multi-modal fusion. To evaluate the superiority of one stream tracker on network complexity, we employ the computational cost to represent the model complexity. We calculate the number of trainable parameters on some two stream trackers, like, and our proposed eMoE-Tracker for the comparison. Results are reported in TABLE VI.
@@ V CONCLUSIONS
Limitations. Despite the superior performance for RGB-E tracking, the limitation of our model is highly dependent on the manual disentanglement annotation for the environmental attributes, thus put restrictions on the generalization of the model. Additionally, the real-world scenarios are complicated and hard to fully consider all the environmental attributes manually. In the future, we expect to learn an agent to obtain the environmental attributes in a learnable manner for multi-modal tracking tasks.
@@ VIII-B Attributes Performance
Attention Map. Since the VisEvent and COESOT dataset all provide 17 attributes for tracking performance evaluation, we can leverage the Matlab toolkit to plot the diagram of curves about the attributes performance comparison. We represent the diagram of curves on 16 attributes except for the no motion scenario in Fig. 11 and Fig. 10, which show the superior performance under all the environmental conditions compared to other SOTA priors.
@@
Figure 9: The visualization of the feature maps and assembling weights for four environmental experts branches under specific scenarios. The 4-digits vector is the attributes label and the following numbers are the corresponding learnable weights for attribute-specific features. The last column denotes the assembled features.
Figure 10: The precision diagram of curves on attributes performance on COESOT dataset. Here illustrates 16 attributes performance, including Background Object Motion, Over Exposure, Illumination Variation, Rotation, Fast Motion, Aspect Ration Change, Motion Blur, Background Clutter, Scale Varation, Viewpoint Change, Partial Occlusion, Out-of-View, Low Illumination, Full Occlusion, Deformation, and Camera Motion.
Figure 11: The precision diagram of curves on attributes performance on VisEvent dataset. Here illustrates 16 attributes performance, including Background Object Motion, Over Exposure, Illumination Variation, Rotation, Fast Motion, Aspect Ration Change, Motion Blur, Background Clutter, Scale Varation, Viewpoint Change, Partial Occlusion, Out-of-View, Low Illumination, Full Occlusion, Deformation, and Camera Motion.
@@ VIII-C Real-World Testing
For real-world testing, we further validate our method on another RGB-event dataset, namely VOT-DVS, to demonstrate the practicality. The VOT-DVS dataset is a popular synthesized RGB-event tracking dataset. The testing result is reported in Table. VII.

### v3

@@ IV-C Ablation Studies
Effectiveness of eMoE and CRM. To validate the effectiveness of our proposed modules, we perform the ablation studies on the VisEvent and COESOT datasets. We implement four comparison experiments inside the network. They are: 1) backbone network 2) backbone network with eMoE; 3) backbone network with eMoE and CRM. The ablation studies can be found in TABLE III. As observed, the best performance happens when we combine all the proposed modules into the backbone network. For the VisEvent dataset, with the incorporation of eMoE and CRM, our method outperforms the backbone model by 7.9% and 6.9% on metrics SR and PR, respectively. Additionally, adding the CRM module to model ①, the SR and PR gain improvements of 2.1% and 0.6%, respectively. The results showcase the effectiveness of the proposed modules.
Analysis on environmental attributes. In the dataset COESOT, there are 17 challenging environmental attributes annotated to help analyze the performance under different challenging conditions. Here we illustrate the overall performance on the COESOT dataset for the disentangled four challenging attributes: illumination variation, motion blur, scale variance, and full occlusion. In Fig. 7, we can find our proposed model eMoE-Tracker outperforms the backbone model and ViPT [10]. Specifically, it achieves 63.6% in occlusion, 77.9% in illumination variance, 73.5% in motion blur, and 80.7% in scale variance on PR, respectively. The results demonstrate that our algorithm is effective in improving tracking precision and robustness under various challenging scenarios.
Inserted layers of eMoE. We achieve RBG-E tracking under various challenging conditions by injecting eMoE blocks into different layers of the backbone model. It is intuitive to investigate the effect on the number of inserted blocks. Here we set different insert intervals for blocks to insert and the intervals are 1,2 4, 6, and 12. Therefore, the first means that all the layers are fully inserted and the last one only inserts the blocks in the last high-level layer. As shown in Fig. V, the best performance can be obtained when the tracking backbone network is fully inserted.
@@
Analysis on the number of experts. Due to the complicated environmental attributes, it is worthwhile to consider the impact of the number of experts on tracking performance. Intuitively, more experts are more powerful at addressing complex environments and decomposing them into environmental attributes for easier tracking. However, on the one hand, too many experts increase the burden on the model and might result in overfitting, on the other hand, the number of experts is unfeasible to extend over four due to the limitation of manual annotations. Therefore, we conduct ablation studies on the number of experts, and the results can be found in TABLE IV.
Analysis on the model complexity. As we mentioned previously, trackers with the two-stream structure, e.g.siamese-based trackers, suffer from model complexity due to the high demand for multi-modal fusion. To evaluate the superiority of one stream tracker on network complexity, we calculate the number of trainable parameters on some two stream trackers, like AFNet [7] and FENet [6], and our proposed eMoE-Tracker for the comparison. Results are reported in TABLE VI.
@@ V CONCLUSIONS
Limitations. Despite the superior performance for RGB-E tracking, the limitation of our model is highly dependent on the manual video-level attribute annotations for the environmental attributes, thus putting restrictions on the generalization of the model. Also, video-level annotations might not provide absolute precise labels for every video sequence in some conditions. In the future, we expect to learn an agent to obtain the environmental attributes in a learnable manner for multi-modal tracking tasks.
@@ VI-A Datasets
Attributes Selection. For the four selected attributes, i.e., illumination variation, motion blur, scale variance and occlusion, on the one hand, they are categorized based on the 17 attributes in the testing sets, on the other hand, they can be based summarized on the observation as some samples shown in Fig. VIII.
@@ VII-B Ablation Studies
Figure 8: RGB image samples from COESOT dataset.
@@ VIII-B Attributes Performance
Attention Map. Since the VisEvent and COESOT datasets all provide 17 attributes for tracking performance evaluation, we can leverage the Matlab toolkit to plot the diagram of curves about the attributes performance comparison. We represent the diagram of curves on 16 attributes except for the no motion scenario in Fig. 12 and Fig. 11, which show the superior performance under all the environmental conditions compared to other SOTA priors.
@@
Figure 10: The visualization of the feature maps and assembling weights for four environmental experts branches under specific scenarios. The 4-digit vector is the attributes label and the following numbers are the corresponding learnable weights for attribute-specific features. The last column denotes the assembled features.
Figure 11: The precision diagram of curves on attributes performance on COESOT dataset. We illustrate performance on 17 attributes, including Background Object Motion, Over-exposure, Illumination Variation, Rotation, Fast Motion, Aspect Ratio Change, Motion Blur, Background Clutter, Scale Variation, Viewpoint Change, Partial Occlusion, Out-of-View, Low Illumination, Full Occlusion, Deformation, and Camera Motion.
Figure 12: The precision diagram of curves on attributes performance on VisEvent dataset. We illustrate performance on 17 attributes, including Background Object Motion, Over-exposure, Illumination Variation, Rotation, Fast Motion, Aspect Ratio Change, Motion Blur, Background Clutter, Scale Variation, Viewpoint Change, Partial Occlusion, Out-of-View, Low Illumination, Full Occlusion, Deformation, and Camera Motion.
@@ VIII-C Real-World Testing
For real-world testing, we further validate our method on another RGB-event dataset, namely FELT-SOT [wang2024FELTSOT], to demonstrate the practicality. The FELT-SOT [wang2024FELTSOT] dataset is a new long-term and large-scale frame-event single object tracking dataset. The testing result is reported in Table. VII.
VIII-D Inference Speed
In object tracking tasks, Frames Per Second (FPS) is used to evaluate the inference speed of tracking methods. For deploying the tracking model on real machines, the inference speed should be greater than 25 FPS. We calculate the inference speed of our model and compare it with existing priors. The result is illustrated in Table. VIII.
TABLE VIII: Comparison of inference speed between our eMoE-Tracker and other existing priors.
Trackers OSTrack [22] ViPT [10] CEUTrack [4] eMoE-Tracker
Inference speed 93.1 49 75 46